OpenStudy (anonymous):

HELP PLEASE What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (anonymous):

first find the slope

OpenStudy (anonymous):

then use one set of points plug them into the y=mx+b equation m=slope then that will give you the b which is the y intecept

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

the slope is y2-y1/x2-x1 so 7-3/2-(-2) =1

OpenStudy (anonymous):

so then it would be y= 1x+b to solve for b plug in (2,7) for x and y therefore 7=2(1)+b b=5

OpenStudy (anonymous):

4/4? which is one?

OpenStudy (anonymous):

yes the answer for the slope is 1 and the variable for slope is m

OpenStudy (anonymous):

the equation after you find b will be y=1x+5 but the answer is with x and y on one side so subtract y from the left and subtract 5 from the right making it -5=x-y

OpenStudy (anonymous):

Alright, thank you.

OpenStudy (anonymous):

yup no problem